### v1

# Counting normals to closed curves in $\mathbb{R}^{3}$

Abstract.
We prove the following results:
(1) For every generic closed smooth curve in $\mathbb{R}^{3}$, there is a point with at least $6$ emanating normals to the curve. If the curve is knotted, there is a point with at least $8$ emanating normals.
(2) For every generic closed piecewise linear curve in $\mathbb{R}^{3}$, there is a point with at least $8$ emanating normals to the curve. If the curve is knotted, there is a point with at least $10$ emanating normals.
The proof is based on the Morse theory for the squared distance function and self intersections of the focal surface.
Key words and phrases:
Space curves, knots, normals. MSC 53A04, 57K10
The second author acknowledges support from the project ‘‘Singularities and Applications’’ - CF 132/31.07.2023 funded by the European Union - NextGenerationEU - through Romania’s National Recovery and Resilience Plan, and support by the grant CNRS-INSMI-IEA-329.
The research was carried on in Batumi during the Winter Math Weeks 2025 and 2026. We thank our host George Khimshiashvili for hospitality and useful remarks.
1. Preliminaries and basic definitions
It has long been conjectured that for every convex body $P\subset\mathbb{R}^{n}$, there exists a point in its interior which belongs to at least $2n$ normals from different points on the boundary of $P$. The conjecture is proven for $n\leq 4$ [3, 8], and almost nothing is known for higher dimensions.
A similar problem, with appropriate definitions, makes sense for convex polytopes. It was proven recently [6, 7] that each convex simple polytope $P\subset\mathbb{R}^{n}$ has a point with at least $2n+4$ normals (provided that $n>2$).
In particular, for curves in the plane, elementary facts are:
For every smooth convex curve in the plane, there is a point with at least four normals to the curve. The bound is exact since an ellipse has no point with more than four normals.
For every piecewise linear convex curve in the plane, there is a point with at least six normals to the curve. This bound is also exact since a triangle has no point with more than six normals.
In the paper we turn to closed curves in $\mathbb{R}^{3}$, either smooth or piecewise linear.
Definition 1.
A segment $yx$ is called a normal to a curve $C$, emanating from the point $y$, if $x\in C$ and $x$ is either a local maximum or a local minimum of the squared distance function
$SQD_{y}:C\rightarrow\mathbb{R},\ \ \ SQD_{y}(x):=|x-y|^{2}.$
The number of normals emanating from $y$ is denoted by $N(y)$.
For a smooth curve $C$, a normal is orthogonal to the tangent vector at the point $x$.
In the piecewise linear case the definition agrees with the geometric intuition: a segment $yx$ is a normal to $C$ if and only if the plane containing $x$ and orthogonal to $xy$ intersects the curve $C$ at the point $x$ non-transversally.
Our main results are:
Theorem 1.
For every generic smooth closed curve in $\mathbb{R}^{3}$, there exists a point $y$ with at least $6$ emanating normals.
If the curve is knotted, there is a point with at least $8$ normals.
As our experience [5, 6, 7] suggests, piecewise linear objects have more normals than smooth ones. So the reader should not be surprised to see the following theorem:
Theorem 2.
For every generic non-planar piecewise linear curve in $\mathbb{R}^{3}$, there is a point with at least $8$ emanating normals.
If the curve is knotted, there is a point with at least $10$ normals.
2. Smooth curves. Proof of Theorem 1
Throughout this section we assume that $C$ is a generic closed smooth curve embedded in $\mathbb{R}^{3}$. From now on, we call local maxima and local minima simply maxima and minima, respectively, for brevity.
A simple observation is: the number of minima equals the number of maxima, and they alternate on the curve (between each two minima there is a maximum, and vice versa).
Lemma 1.
For an embedded closed curve $C$, either smooth or piecewise linear, if there exists $y$ with $N(y)=2$, then $C$ is an unknot.
Proof.
We mimic one of the proofs of the Milnor-Fáry theorem, see [9]. Consider the family of concentric spheres $S_{r}$ of radius $r$ centered at $y$. By assumption, each such sphere intersects $C$ at most twice. By a slight perturbation of $C$, we may assume that the two intersection points are never antipodal. Connect the intersection points by the (uniquely defined) shortest geodesic segment. The union of all such segments is an embedded disk $d$ with the property $C=\partial d$, hence $C$ is unknotted. ∎
For $x\in C$ let us fix the following notation: $h(x)$ is the osculating plane to $C$ at the point $x$, the point $F(x)\in h(x)$ is the center of the osculating circle. A generic curve has an everywhere non-zero curvature, so we always assume that $F(x)$ is not at the infinity.
$L(x)$ is the line passing through $F(x)$ and orthogonal to $h(x)$. That is, $L(x)$ goes in the direction of the binormal $b(x)=t(x)\times n(x)$. The focal surface $\mathcal{F}(C)$ is the union of all these lines. We refer the reader to [2] and [10] for more details on spatial curves.
We recall that $\mathcal{F}(C)$ is the bifurcation set of the function $SQD_{y}$. That is, $\mathcal{F}(C)$ cuts the space into chambers. In each of the chambers the number of normals $N(y)$ persists, and if $y$ crosses a sheet of $\mathcal{F}(C)$ transversally, either two normals are born, or two normals die.
Proposition 1.
The focal surface $\mathcal{F}(C)$ of any generic closed smooth curve in $\mathbb{R}^{3}$ has a transverse double self-intersection.
Proof.
The focal surface $\mathcal{F}(C)$ is equal to the tangent developable of the evolute $\mathcal{E}(C)$, that is, $\mathcal{F}(C)$ is the union of all tangent lines to $\mathcal{E}(C)$ [2]. The evolute is generically a closed curve, possibly with cusps and points at infinity.
  1. (1)
    Assume that the torsion of $C$ never vanishes. Then $\mathcal{E}(C)$ has no points at infinity. Consider the following two cases:
    (a) $\mathcal{E}(C)$ has no cusps. We have to prove that that $\mathcal{E}(C)$ has two intersecting tangent lines. Take a support plane to the curve $\mathcal{E}$. It is a plane that is tangential at some point $x\in\mathcal{E}$ and intersects $C$ at no other point. Start rotating the plane around the tangent line to $\mathcal{E}(C)$ at the point $x$. Sooner or later the plane hits $\mathcal{E}(C)$ at some other point $x^{\prime}$. Both tangent lines (to $x$ and to $x^{\prime}$) lie in the plane and therefore generically intersect.
    (b) Assume that $\mathcal{E}(C)$ has a cusp. The cusp corresponds to a swallow tail of $\mathcal{F}(C)$, which provides a curve of transverse self-intersection.
  2. (2)
    Assume that the torsion of $C$ changes the sign at some point $x_{0}\in C$. We shall prove that there is a pair of intersecting lines $L(x)$ and $L(x^{\prime})$ for some $x$ and $x^{\prime}$ lying near $x_{0}$ (in fact, we show that there is a one-parametric family of such pairs).
    An ordered pair of oriented lines in $\mathbb{R}^{3}$ is skew if they neither intersect nor are parallel. Each skew pair of lines has a naturally defined orientation: we have an ordered triple of vectors, the defining vectors of the lines, and the vector associated to their common perpendicular which is directed from the first line to the second. Then, by definition, the orientation of the pair of lines is the orientation of the triple of vectors. The orientation clearly persists under continuous deformations of the pair provided that the lines remain skew during the deformation.
    For $x\in C$ let us orient the line $L(x)$ the direction of the binormal. Pick two different points $x,x^{\prime}\in C$ close to $x_{0}$ and lying on one side to $x_{0}$. The orientation of the associated pair $L(x),L(x^{\prime})$ depends on the sign of the torsion. Let the pair $x,x^{\prime}$ move along the curve to some target pair of points on the other side with respect to $x_{0}$. Since the orientation of the pair $L(x),L(x^{\prime})$ changes, somewhere in between the lines intersect. This corresponds to self-crossing of $\mathcal{F}(C)$.
Now we are ready to prove Theorem 1:
Proof.
A neighborhood of a point of transverse self-intersection of $\mathcal{F}(C)$ is divided into four chambers by two sheets of $\mathcal{F}(C)$. Crossing of each sheet changes $N(y)$ by $2$. Each sheet can be given a co-orientation in the direction of the increasing number of normals. It follows, that in these four chambers we have $N$, $N+2$ and $N+4$ emanating normals for some $N$. The value of $N$ is at least $2$, whereas in the knotted case by Lemma 1 it is at least $4$ which implies the desired. ∎
We present below a sketch of yet another proof of the first statement of Theorem 1, based on a different idea.
Proof.
Assuming the contrary, pick a triple of points $a,b,c\in C$ such that the normal planes $a^{\perp},b^{\perp},c^{\perp}$ intersect at a unique point $y=y(a,b,c)$. Then generically $a,b,c$ are some minima and maxima of $SQD_{y}$. If they are all minima (resp., all maxima), then there are necessarily also $3$ maxima (resp., $3$ minima), and hence $N(y)\geq 6$.
So assume that $a,b,c$ yields (max,max,min). There is exactly one more minimum of $SQD_{y}$, otherwise there are $6$ normals. Let the triple $a_{t},b_{t},c_{t}$ move continuously along the curve from $a,b,c$ to $b,c,a$. Eventually we get (max,min,max). We may assume that the path is such that the point $y(a_{t},b_{t},c_{t})$ intersects the bifurcation set $\mathcal{F}(C)$ transversally. So we move from (max,max,min) to (max,min,max). Therefore on the way there is either a bifurcation, or $y(a_{t},b_{t},c_{t})$ passes through infinity. Let us analyze these two possibilities.
(1) A bifurcation means either birth or death of two critical points. If two new critical points arise, we have six of them right after the bifurcation. In our setting the death of two critical points is impossible since on the path we have always at least four of them. So there are no bifurcations.
(2) Passing through infinity turns minima to maxima and vice versa. So (max,max,min) may turn to (min,min,max) but not to (max,min,max).
Example 1.
A planar ellipse in $\mathbb{R}^{3}$ has no point with $6$ normals. However, its small perturbation yields points with $6$ normals.
A natural question is about the maximum of $N(y)$. Clearly, inside the isotopy class of the curve one can arrange arbitrary high numbers. If one restricts to algebraic curves there will be some upper bounds. This bound is especially studied in the context of Euclidean Distance degree (ED-degree) [1]. One considers as upper bound the number of critical points of the complexification of the variety with respect to the complex squared distance function. This number does not depend on $y$.
For smooth complete intersection curves in $\mathbb{R}^{3}$ with degree $(d_{1},d_{2})$ the ED-degree is $d_{1}d_{2}(d_{1}+d_{2}-2)$.
In the case of the lifted ellipse one could take $d_{1}=d_{2}=2$, so the bound is $8$.
3. Piecewise linear curves. Proof of Theorem 2
In this section we assume that $C$ is a piecewise linear non-planar generic curve embedded in $\mathbb{R}^{3}$. In this case the maxima of $SQD_{y}$ are always attained at the vertices, whereas minima may be attained at both vertices and edges.
We start with the proof of the first statement of Theorem 2:
Proof.
There exists a superscribed sphere, that is, a sphere containing $4$ vertices of the curve. Take its center $O$. The function $SQD_{O}$ has at least four maxima, hence four minima. ∎
Now let us consider a knotted curve $C$ and prove the second statement.
Proof.
Let us first describe the bifurcation set $\mathcal{B}(C)$. Each vertex $v$ of the curve with two emanating edges $l$ and $l^{\prime}$ yields two half-planes with a common boundary. This is how it arises: take normal planes $h$ and $h^{\prime}$ to the edges $l$ and $l^{\prime}$ such that $v\in l,l^{\prime}$. Taken together, they divide the space into four domains. In exactly one of these domains the vertex $v$ is a maximum of $SQD$. The union of the two half-planes is the boundary of this domain; we call it the wedge related to the vertex $v$. The intersection $h\cap h^{\prime}$ is called the ridge of the wedge, see Fig. 1.
Figure 1. Each vertex of $C$ gives a wedge of $\mathcal{B}(C)$.
Each wedge cuts the space $\mathbb{R}^{3}$ into two connected components. If $y$ belongs to the smaller one, $yv$ is a normal which corresponds to a maximum.
Whenever a point $y$ crosses a wedge transversely, there is a bifurcation involving the vertex $v$. This means that two normals (one maximum related to $v$ and the other minimum related to either $l$ or $l^{\prime}$) either die of is born.
By definition, the union of all the wedges equals $\mathcal{B}(C)$. The set $\mathcal{B}(C)$ cuts the space $\mathbb{R}^{3}$ into chambers. Each chamber is some polytope, possibly non-convex, possibly unbounded. If a point moves in one and the same chamber (that is, crosses no wedge), the number of normals persists.
Proposition 2.
There is a point of triple intersection of three different wedges of $\mathcal{B}(C)$.
Let us first show that the proposition implies the theorem. A point of triple intersection has $8$ incident chambers. The number of normals for these chambers is N,N+2,N+4,N,\ N+2,\ N+4, and $N+6$ for some $N$. By Lemma 1, we have $N\geq 4$, which completes the proof.
Now let us prove the proposition.
Given a chamber, some of its edges are intersections of two different wedges, whereas some edges come from ridges. Call the latter ones ridge edges.
Genericity implies:
Lemma 2.
  1. (1)
    No chamber contains a straight line.
  2. (2)
    A face of a chamber has no more than one ridge edge.
  3. (3)
    Each vertex of a chamber has three emanating edges of the chamber. At most one of them is a ridge edge.
Proof.
The statements (2) and (3) follow from genericity, so let us prove (1).
It suffices to prove that each line $L$ intersects the bifurcation set. Take the vertex $v$ of $C$ such that the function $SQD_{y}$ attains a local maximum at $v$ for some point $y\in L$. If $L$ does not cross $\mathcal{B}(C)$, the function $SQD_{y}$ attains maximum at $v$ for all $y\in L$. This happens only if $L$ is parallel to a ridge associated to the vertex $v$. On the other hand, in the knotted case such a vertex $v$ is not unique by Lemma 1. By genericity, no two ridges are parallel, so we have a contradiction. ∎
Lemma 3.
If a bounded face of a chamber is not a triangle, then it has a vertex with no incident ridge edge. Such a vertex is a transverse intersection of three wedges.
Proof.
Let the face be a quadrilateral $ABCD$ (the bigger numbers of vertices are treated similarly).
Assuming the contrary (each vertex has an emanating ridge edge), consider two cases:
(1) None of the edges of $ABCD$ is a ridge. These points have one more incident edge of the chamber. All of them are ridges, which contradicts Lemma 3, (2).
(2) Assume $AB$ is a ridge. Then either $C$ or $D$ has no incident ridge edges by Lemma 3, (2).
So it remains to show that not all the bounded faces of $\mathcal{B}(C)$ are triangles.
Consider the sphere at infinity. Each wedge intersects it by a lune which is the union of two geodesics connecting the pair of antipodal points that correspond to the ridge of the wedge. There is a natural circle ordering on the lunes since they come from vertices on the curve. Two neighbour lunes necessarily have a common geodesic segment, see Fig. 2. If the lunes are not neighbors, we may assume by genericity that they intersect transversally.
Figure 2. Lunes on the sphere at infinity: two lunes, and some different positions of three neighbour lunes.
The union of lunes yields a tiling $T$ of the sphere into some spherical polygons. A simple analysis shows that the vertices of the tiling have valence $4$ with the following possible exception at the endpoints of the lunes:
Lemma 4.
Each lune has either has two endpoints of valency $3$, or one endpoint of valency $2$ and the other endpoint of valency $4$. All the other vertices of $T$ have valency $4$.
Proof.
See Fig. 2 for some possible relative positions of lunes. The reader can easily fill in all the missing cases. ∎
Lemma 5.
Each tile of $T$ fits in an open hemisphere. Therefore, no tile is a bigon (no tile is a lune).
Proof.
By Lemma 1, each point of the sphere (and therefore, each tile) is covered by at least two lunes. ∎
Lemma 6.
Not all the tiles of $T$ are triangles and quadrilaterals.
Proof.
We use the following common approach: assuming the contrary, replace each triangular tile by an equilateral triangle with unit edge lengths, and each quadrilateral by a square also with the unit edge lengths. They patch together to a sphere with almost everywhere flat metric. The curvature is concentrated at the vertices, and each vertex contributes the angular defect (it equals the difference of $2\pi$ and the sum of the incident angles). The total curvature can be estimated as follows: each vertex of valency $4$ contributes at least zero; each vertex of valency $3$ contributes at least $\pi/2$, and each vertex of valency $2$ contributes at least $\pi$. Since the sum of the curvatures over all the vertices equals $4\pi$, we conclude that there are at most $4$ lunes. Since the number of lunes equals the number of vertices of $C$, we have a contradiction. ∎
We proceed with the proof of the proposition. So the tiling $T$ has a tile $\tau$ with at least $5$ vertices. Let us assume that there are exactly $5$ vertices; the bigger numbers are treated similarly. There is an unbounded chamber $\sigma$ adjacent to this tile. All the vertices of $\sigma$ are trivalent because of genericity.
Figure 3. The chamber $\sigma$ (left). The proof does not work for quadrilateral $\tau$ (right).
By similar curvature count we conclude that $\sigma$ has a bounded face with at least $4$ vertices.
First observe that all the points $A^{\prime},B^{\prime},C^{\prime},D^{\prime},E^{\prime}$ are distinct (see Fig. 3).
Next, replace the tile $\tau$ by a regular pentagon $ABCDE$, replace all other faces by regular polygons with equal edge lengths. Assume that all the bounded faces are triangles. Then the contribution of $A,B,C,D,E$ equals at least
$5\cdot(2\pi-(\pi/2+\pi/2+{3}\pi/{5}))=2\pi.$
The contribution of $A^{\prime},B^{\prime},C^{\prime},D^{\prime},E^{\prime}$ is at least
$5\cdot(2\pi-(\pi+\pi/3))=\frac{10\pi}{3},$
which is already bigger than $4\pi$. [^1] The genus of the chamber might be non-zero, but in this case we come to a contradiction as well.
Remark. For a quadrilateral $\tau$ the proof breaks down: the points $A^{\prime},B^{\prime},C^{\prime},D^{\prime}$ might be not all distinct, see Fig. 3, right.
The lemma (and hence the proposition) are proven. ∎
As a by-product of the above, we get the following.
Proposition 3.
If $C$ is a knotted piecewise linear curve, then there exists a linear function whose restriction on $C$ has at least $8$ critical points (that is, minima and maxima).
Proof.
There are points of transverse intersections of two different sheets on the infinity. Since for every chamber, the number of normals is at least $4$, and crossing of a sheet amounts to death (or birth) of yet another maximum and minimum, we get the desired. ∎
Remark. For closed PL-curves the number of normals is always finite and one has $N(y)\leq 2v$, where $v$ is the number of vertices.
For any $n$ there are polygons, such that $N(y)=2v$. The concept of aED-degree is defined by the average of $N(y)$ with respect to a measure, cf [1]. So aED-degree $\leq 2v$.